OpenStudy (anonymous):

Can someone show me how 2 simplify (2 √5 +3 √7)² step by step

OpenStudy (anonymous):

so you know that (a+b)^2 = a^2 + 2ab + b^2 ?

OpenStudy (anonymous):

Yeah but i dont know what 2 do with that....

OpenStudy (anonymous):

ahha, you have to use it? So what is \[(2\sqrt{5})^2\]

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Yes. and what is \[(3\sqrt{7)}) ^2\]

OpenStudy (anonymous):

63

OpenStudy (anonymous):

Yes. and what would approx be 2 x \[2\sqrt{5} . 3\sqrt{7}\]

OpenStudy (anonymous):

i mean 2 . 2 root 5 . 3 root 7

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i mean \[(2\sqrt{5)}^2 = 20, (3\sqrt{7)}^2 = 63, what is 2 . 2\sqrt{5} . 3\sqrt{7} ?\]

OpenStudy (anonymous):

166